We wish to know if we can conclude that the mean daily caloric intake in


the adult rural population of


a developing country is less than 2000. A sample of 500 had a mean of 1985


and a standard deviation


of 210. Let alpha = 0.5

1
Expert's answer
2022-03-23T14:52:00-0400

"H_0:\mu \geqslant 2000\\H_1:\mu <2000\\T=\sqrt{n}\frac{\bar{x}-\mu}{s}=\sqrt{500}\frac{1985-2000}{210}=-1.59719\sim t_{n-1}=t_{499}\\P-value:\\P\left( T\leqslant -1.59719 \right) =F_{t,499}\left( -1.59719 \right) =0.0554"

Since the P-value is less than the significance level, the null hypothesis is declined. The mean value is less than 2000.